A permanganate solution is standardized by titrating it with a 0.1049 M Fe2+ solution. The balanced net ionic equation for the reaction is:
MnO4-(aq) + 5Fe2+(aq)+8H3O+(aq)Mn2+(aq) + 5Fe3+(aq)+12H2O(l)

If 22.95 mL of the 0.1049 M Fe2+ solution is required to react completely with 30.00 mL of the permanganate solution, calculate the concentration of the permanganate solution.

__________ M
1
Expert's answer
2019-03-21T06:24:55-0400

"mol(Fe2+)=22.95mL*\dfrac{0.1049mol}{1000mL}=0.002407mol"


"mol(MnO4-)=0.002407mol*\dfrac{1 mol(MnO4-) }{5mol(Fe2+)}=0.0004814mol"

"c(MnO4-)=\dfrac{0.0004814mol}{0.03L}=0.01605 mol/L"